#619 closed flight processing (fixed)

ES17/136, flight day 167b/2017, Albacete

Reported by: asm Owned by:
Priority: immediate Milestone:
Component: Processing: general Keywords:
Cc: Other processors:

Description

Data location: ~arsf/arsf_data/2017/flight_data/spain/ES17_136-2017_167b_Albacete

Data arrived from NERC-ARF via hard drive 2017-06-30

Scientific objective: Develop LLT/SSE using AisaOwl data
Priority: Urgent

PI: Martin Wooster

Sensors:

  • Fenix (flown, requested)
  • Leica LIDAR (flown, not requested)
  • OWL (flown, requested)
  • PhaseOne (flown, requested)

Change History (33)

comment:1 Changed 22 months ago by asm

Unpacking

Project has been unpacked. There are 6 fenix flightlines and 6 owl lines.

Checks have been run:
-> For 1 fenix file, fps differs from fps_qpf by 0.51
-> An owl line has 1 dropped frame
-> quite a few owl files have between 1 and 3 dropped frames incidents
-> 2 lidar files are empty: 170616_142515/LDR170616_142515_0000000005.scn and 170616_143321/LDR170616_143321_0000000006.scn

comment:2 Changed 21 months ago by dac

dGPS data collected at the same time as flight provided by PI and copied to: /users/rsg/arsf/arsf_data/2017/misc/basestation_data/ES17_136_Basestation_Data

comment:3 Changed 21 months ago by dac

Digital Camera

Converted raw images to tiff.

comment:4 Changed 21 months ago by lah

Owl

Processed to level 1b to compare temperature with 167a and 167c.

comment:5 Changed 21 months ago by lah

Logsheet landing time of 17:03 appears incorrect (takes off again at 16:30 for 167c). Doesn't appear to cause a problem in nav processing.

comment:6 Changed 21 months ago by lah

Navigation - Basestation

Lat: 39 03 37.37019
Lon: -2 05 53.41236
Height: 748.751 m

Estimated basestation height of 1.60 m has been used since there is no value in the file.

comment:7 Changed 21 months ago by lah

Navigation Processing

Best position separation obtain with default configuration settings. Keeping the transit part of the flight gives a better result within the flight lines, though does go outside 0.05 at the end of the flight (beyond the data lines).

comment:8 Changed 21 months ago by lah

Owl Processing

Scts still look good at either 2.00 or 3.00. Some slight offsets on part of flight lines, but other parts line up very well.

Flightline Owl SCT
1 3
2 2
3 2
4 2
5 2
6 2

Proceeding to map files.

comment:9 Changed 21 months ago by lah

Owl Processing

Delivery made - Ready for delivery check.

comment:10 Changed 21 months ago by dac

Fenix Processing

Started hyperspectral processing.

comment:11 Changed 21 months ago by dac

Fenix

Found SCT values:

Flightline FENIX
1 2
2 2
3 2
4 2
5 2
6 1

comment:12 Changed 21 months ago by dac

Fenix Processing

Generated delivery. XML files still have ARSF listed in them so changed using:

sed -i "s/ARSF/NERC-ARF/g" *xml
sed -i "s#https://arsf-dan.nerc.ac.uk#https://nerc-arf-dan.pml.ac.uk#g" *xml

Will submit merge request with fix to script.

Ready for delivery check.

Last edited 21 months ago by dac (previous) (diff)

comment:13 Changed 20 months ago by dac

Basestation

NERC GEF retrieved basestation height using Leica GeoOffice. Total height 1.328 m. Will reprocess navigation data with correct height.

comment:14 Changed 20 months ago by dac

Reprocessed navigation data using correct basestation height.

Digital Camera

Tagged images and generated thumbnails.

comment:15 Changed 20 months ago by dac

Hyperspectral and Thermal

Started reprocessing with new .sol file and correct FOV vectors for Fenix.

comment:16 Changed 20 months ago by dac

Hyperspectral

Generated delivery. Ready for checking.

comment:17 Changed 20 months ago by lah

Owl reprocessing

Looks like Owl was sent to reprocess without the level 1b files, so files were not mapped. Have sent level 1b files to reprocess as it's easier than renaming the delivery files.

comment:18 Changed 20 months ago by dac

Hyperspectral

Started reprocessing with bad pixel file.

comment:19 Changed 20 months ago by dac

Hyperspectral

Generated delivery - ready for checking.

comment:20 Changed 20 months ago by lah

Fenix Delivery Check

  • removed fodis directory
  • removed ARF as aircraft base from logsheet
  • checked FOV are now for the loan system
  • proj_tidy doesn't like any of the names, probably because the project name has too many numbers in the second part (136)
  • All line over and underflows look ok
  • Spectra look fine in ENVI
  • Most of the bad pixels have been correctly masked. A few remain in the noisy water absorption bands, but these shouldn't be used anyway.

Marking as ready to deliver and zipping files.

comment:21 Changed 20 months ago by lah

Owl Ready for Delivery Check

Created new Owl delivery. Dornier picture still in readme. Do we have a twin otter picture to replace it with?

comment:22 Changed 20 months ago by dac

Do we have a twin otter picture to replace it with?

Not at the moment. Will remove for now and see if we can get a similar picture of the Twin Otter in flight to use for next season.

comment:23 Changed 20 months ago by dac

Owl delivery check

  • Updated operator in XML from 'Johnson' to 'James Johnson'
  • check_apl_cmd ran through without any problems

Started zipping mapped files. Marking as ready to deliver.

comment:24 Changed 20 months ago by dac

Data delivery

Fenix and Owl data placed on FTP server (slot 4). Notification sent to PI.

comment:25 Changed 18 months ago by asm

Digital Camera

Created delivery and readme file. Only waiting to have the data quality reports to mark the project as awaiting delivery check.

comment:26 Changed 18 months ago by asm

Digital Camera

Added data quality report. Project ready for Delivery Check.

comment:27 Changed 18 months ago by dac

Digital Camera

Starting delivery check.

comment:28 Changed 18 months ago by dac

Digital Camera Delivery Check

All looks good. LiDAR lines in KML didn't look right. There wasn't LiDAR for this flight anyway so removed them.

Ready to deliver.

comment:29 Changed 18 months ago by asm

Camera Delivery

Placed on FTP server (slot 4). Notification sent to PI on 18/10/2017.

comment:30 Changed 17 months ago by asm

Archiving

Tidying up before archive. Will delete Owl temperature directory as the owl sct problem was fixed and files can be recreated.

comment:31 Changed 17 months ago by asm

Archiving

Started uploading data to CEDA.

comment:32 Changed 17 months ago by asm

Archiving

Paused the upload of data to CEDA.

comment:33 Changed 17 months ago by dac

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.